OpenStudy (anonymous):

how would you solve sin^2x + cosx + 1?

OpenStudy (anonymous):

Could we rewrite sin^2x

OpenStudy (turingtest):

use the identity\[\sin^2x+\cos^2x=1\]and sub in for \(\sin^2x\)

OpenStudy (anonymous):

so (1 - cos^2x) + 1

OpenStudy (anonymous):

then combine the 1s?

OpenStudy (anonymous):

yeah plus you got that other cosine still there, so you end up with a quadrtic equation (in cosine)

OpenStudy (anonymous):

\[\sin^2(x) + \cos(x) + 1\]hold on just noticed there was no equal sign

OpenStudy (anonymous):

is it \[\sin^2(x) + \cos(x) + 1=0\]?

OpenStudy (anonymous):

(1 - cos^2x) + cosx + 1, the goal is to simplify I think?

OpenStudy (anonymous):

wait, YES!

OpenStudy (anonymous):

i don't know why that would be any simpler

OpenStudy (anonymous):

it's = 0 @satellite73

OpenStudy (anonymous):

whew not it make sense

OpenStudy (anonymous):

\[\sin^2(x) + \cos(x) + 1=0\] \[(1-\cos^2(x)+\cos(x)+1=0\] \[-\cos^2(x)+\cos(x)+2=0\] \[\cos^2(x)-\cos(x)-2=0\] solve as you would \[z^2-z-2=0\] by factoring

OpenStudy (anonymous):

thank you so much! we have a test tomorrow, but for some reason by looking at the equations, i've been overcomplicating everything, and forgetting how to solve...

OpenStudy (anonymous):

when factoring, don't I have to check for extraneous solutions?

OpenStudy (anonymous):

\[(\cos(x)-2)(\cos(x)+1)=0\] \[\cos(x)=2\] which you can ignore because it never is 2 and \[\cos(x)=-1\] which you can solve easily

OpenStudy (anonymous):

extranous or whatever your job is to solve for \(x\) not \(\cos(x)\) so you still have to complete the problem, not hard in this case since \(\cos(x)=-1\) at \(x=\pi, 3\pi, 5\pi, ...\)

OpenStudy (anonymous):

good luck on your exam!!

OpenStudy (anonymous):

alright, thank you so much! you've helped so much!